# Exploring the cosmological degeneracy between decaying dark matter model and viscous $\Lambda$CDM

### v1

@@ II.3 Modified $\Lambda$CDM model
Notice that the first term in the previous expression resembles the Eckart model, $\Pi=-3H_{0}E_{\Lambda}(z)\xi(z)$ with viscosity coefficient, $\xi(z)=H_{0}E_{\Lambda}(z)$ eckart . However, the above result indicates that we are dealing with a more general description for viscous effects and our construction does not depend on the usual and intricate constitutive equations which describe such effects; these equations depend on the framework used to introduce the viscosity effects in the cosmic fluid. According to the form of Eq. (24) our construction is more general than the one considered in velten , where the Eckart model was proved. It is worthy to mention that in this case the viscous pressure is reconstructed and characterized by the physical parameters $\alpha$, $\Omega_{i}$ and $H_{0}$ coming from the decaying dark matter model instead of the usual arbitrary parameters $\xi_{0}$ and $s$. In general, the election of the values for the parameters $\xi_{0}$ and $s$ is only justified and motivated by the mathematical simplification that can be obtained on the dynamical equation obeyed by the Hubble parameter in the viscous description.
@@
In Fig. (2) we show the behavior of the viscous pressure (24) as function of the red-shift for different values of the decaying rate. In any case we observe that $\Pi$ suffers a transition to an attractor behavior around $z\sim 0.25$. After this point (for $z\leq 0.25$) and all along the future cosmic history, $\Pi$ evolves exactly in the same way independently of the value of the decay rate $\alpha$. At the far future, the viscosity pressure in the phantom scenario tends to become more negative as the universe evolves, whilst for the quintessence case it remains negative but close to zero. At times previous to the attractor transition, the behavior depends on the value of $\alpha$. Namely, for very small $\alpha$ values, $\Pi$ takes large values at the far past and decreases along the cosmic evolution either for quintessence and phantom dark energy until today. In contrast, for larger values of the decaying rate ($\alpha>0.1$) in both cases, $\Pi$ increases from negative values in the distant past, towards positive ones. For the phantom case, this transition occurs before than for the quintessence case. For $0.1<\alpha\leq 0.6$ positive phase of $\Pi$ in the past lasts a short while and finishes around the attractor transition. In the extreme case with $\alpha>0.8$ this phase does not occur. In addition, we computed the deceleration parameter (See figure (3)) and we observe that it exhibits a transition from positive to negative before the change of sign in $\Pi$. On the other hand, as the value of the parameter $\alpha$ increases such transition for $q$ is lost, the deceleration parameter is always negative. This is interesting as the viscous effects give rise to an accelerated expansion even for the quintessence case in which the state parameter of dark energy is not as negative as for a CC. Notice that since in our formulation we have $\Lambda\rightarrow\Lambda(z)$, the behavior of $q$ at late times deviates from the expected value $-1$ obtained for the CC due to the existence of the viscous contribution in the matter sector. However, as can be seen in the plot, such attractor can be recovered by considering the adequate value for $\omega_{\mathrm{de}}$.
@@
Finally, using the Eq. (8) with $p=0$ we can construct the $\Gamma$ term if we interpret $\Pi$ as the creation pressure, in this case we have
@@
where $\Omega_{\mathrm{vm}}(z)$ is obtained from Eq. (22) and $\Pi$ can be taken from (24). The source term obtained in this description deviates from the usual functional form used in the literature for $\Gamma$ gamma , i.e., $\Gamma\propto E^{m}(z)$ being $m$ a constant. In this case we observe, $\Gamma=\Gamma(E,E^{\prime})$, where the prime denotes derivatives w.r.t. red-shift. This behavior is interesting since incorporates matter couplings with curvature, see for instance Ref. odintsov . Given the form of $\Pi$, we observe that the particle production behaves in a different way for quintessence and phantom dark energy and will be well defined always that $\Gamma>0$ since the particle production is possible due to the transfer of energy from the gravitational field to matter creation .
@@
III Testing the model against observations
III.1 Test of the expansion of the universe with Supernovae AI from Union 2.2
In this section we aim to test whether our model provides a good description of the accelerated expansion of the universe. We aim to compare our theoretical predictions of distance modulus as function of red-shift with measurements for the sample of supernovae type IA from the Supernovae Cosmology Project compilation Union 2.2 Suzuki:2012 . Firstly, we computed the theoretical predictions for the luminous distance as function of red-shift given as:
@@ III Testing the model against observations
Where $c$ is the speed of light and the denominator in the integrand corresponds to that given by (23). We carried out the corresponding numerical integrations by implementing the trapezoid method in our own python code. Since the truncation error of this method is order 2 of the step size, then the accuracy of our results is good enough for our purposes. Afterwards we computed the modulus simply by using:
@@
Figure 4: Predicted distance modulus as function of red-shift from our model (purple line) overlapped to the observed modulus for Union 2.1 supernovae. The values of the parameters for this model corresponds to $\Omega_{m}h^{2}=0.126$, $H_{0}=73\,\mathrm{km/s/Mpc}$, $\omega_{0}=-0.7$ and $\alpha=0.5$.
As shown in figure (4), as the decay rate parameter $\alpha$ increases a larger distance modulus is expected and such modification is more pronounced at large red-shift, while predictions for small red-shifts remain indistinguishable from those from $\Lambda$CDM. Given the small sensitivity of $\mu$ to variations of $\alpha$ it is expected that this data is no able to strongly constrain this parameter.
@@ III.2 Test of the expansion of the universe with BAO
In order to strengthen our constrains from SNaI distance measurements, we additionally consider into our analysis measurements of the BAO feature along the line-of-sight and transverse directions can separately measure $H(z)$ and the comoving angular diameter distance $D_{M}(z)$ from SDSS-III data Alam:2017 . Variations in the cosmological parameters or the pre-recombination energy density can alter the sound horizon of acoustic oscillations given by:
@@
where $\mathcal{R}=\frac{3\rho_{b}(z)}{4\rho_{\gamma}(z)}=\frac{3\Omega_{b0}}{4\Omega_{\gamma 0}}(z+1)^{-1}$ and $z^{*}$ corresponds to the red-shift at recombination. Therefore, measurements of BAO actually constrain the combinations of $D_{M}(z)/r_{s}$ and $H(z)r_{s}$, therefore we compare predictions of $D_{M}(z)r_{s,fid}/r_{s}$ and $H(z)r_{s}/r_{s,fid}$ with their inferred values from observations of BAO by BOSS-DR12, where $r_{s,fid}$ corresponds to the $\Lambda$CDM fiducial value of the sound horizon given as $r_{s,fid}=147.78\,\mathrm{Mpc}$ which is used as normalization constant. In Figure (5) predictions of $D_{M}(z)r_{s,fid}/r_{s}$ are shown for different models corresponding to a range of values of $\alpha$. Clearly, this quantity barely depends on $\alpha$ at low red-shift, on the contrary, it is sensitive to this parameter at high red-shift. Within this range $\alpha\sim 0.5$ seems to fit properly the observational data. On the other hand, an angle-averaged galaxy BAO measurement constrains the following combination Alam:2017 :
@@ III.3 Bayesian Analysis
In order to obtain baseline estimates for the parameters of the viscous $\Lambda$CDM model, we sample the corresponding parameter space by using the Markov-Chain Monte-Carlo (MCMC) method. Specifically, we implemented the Metropolis-Hasting algorithm in our own python code in order construct a set of Markov-chains which serve as samples in the space of parameters of our model used to determine the posterior distribution function corresponding to supernovae data from Union 2.1 and BAO data from BOSS-DR12 Suzuki:2012 ; Alam:2017 . We marginalized the posterior distribution in order to derive estimations of the best-fit parameters of our model and their uncertainty at $68\%$ and $95\%$ level of confidence (See Figure (6) and the table at the end of this section).
@@
Interestingly, the marginalized distribution for $\alpha$ is pretty similar according either to SNaI and BAO, with a best-fit value $\alpha_{bf}\simeq 0.5$. The fact that $\alpha\neq 0$ gives rise to a better fit of both datasets suggests that the occurrence of dissipative effects along the expansion history is not only feasible but rather, the dynamics of expansion might be better described in comparison to the standard $\Lambda$CDM. Another point that is important to stress is that marginalized distributions either for $\Omega_{m}h^{2}$ and $H_{0}$ are consistent even at $1-\sigma$ for both datasets, SNaI and BAO. A slight tension can be noticed for the dark energy state parameter $\omega_{0}$. Although of this discrepancy in the estimate of $\omega_{0}$ for both datasets is not statistically significant, it can be due to the different features of the expansion of the universe at late and early times. Also, it is worth noticing that estimates for these parameters according to BAO are stronger than those for SNaI, this can be explained by the large sensitivity of the theoretical function $D_{M}(z)$ to these parameters at the largest red-shift values in the sample and to the smallness of the error bars of BAO measurements (See figure (5)).
IV Final Remarks
In this work we study a cosmological model in which a generic candidate of dark matter undergoes into a exponential decay into relativistic dark particles at the late stages of evolution of the universe. Since the decay process of dark matter occurs out of chemical equilibrium given that the average kinetic energy is lower than the mass of the candidate, the entropy grows considerably and therefore it is irreversible. As a consequence, the fluid turns out to be not perfect and a bulk viscosity pressure can be associated to it.
On the other hand it is well know that different cosmologies describe identically the background evolution of the universe, that is, the solutions for the scale factor are degenerated between models. Therefore, by virtue of this degeneracy at background level a mapping between the space of parameters of different models can be established. In this work we map the parameters of the cosmological model including decaying dark matter described above and a slight extension of the standard cosmological model in which dark matter corresponds to an effective viscous fluid and dark energy corresponding to a perfect fluid with barotropic equation of state. By setting this mapping we are able to compute an estimate of the bulk viscosity of decaying dark matter in an effective manner and to describe the expansion of the universe according to this framework. As discussed previously, the reconstructed viscous pressure evolves in a different way depending on the value of the parameter $\omega_{\mathrm{de}}$. However, for phantom dark energy and quintessence, it has a negative contribution to the pressure of the dark matter sector from the recent past ($z\leq 1$) until the far future ($z=-1$). If we consider the same region for the deceleration parameter we can observe a transition from decelerated to accelerated cosmic expansion for $\alpha\leq 0.6$. The behavior at the far future for this parameter deviates from the one expected in the CC scenario since in our case due to the mapping between models, we must reinterpret the CC as a variable dark energy. It is worthy to mention that despite the consideration of the phantom dark energy, $\omega_{\mathrm{de}}<-1$; the model does not exhibit a true (big rip) singularity, in this case the singularity is allowed until infinite time has elapsed, this is known as little rip and it has been explored previously within viscous cosmologies.
In addition, by using Bayesian statistics, we test the expansion of the universe predicted by our model against observations of luminous distance and red-shift of type IA supernovae from the Union 2.2 sample and distance measurements from BAO detected in SDSS catalogs. By sampling the space of parameters of our model using the Markov-Chain-Monte-Carlo method implemented by means of the Metropolis-Hastings algorithm in our own Python code, we derive the corresponding 1D and 2D posterior distributions from which we obtain estimates of the model parameters and their uncertainties. As main result of this analysis we found that dissipative effects along the cosmic history are feasible and could be able to provide a better description of the cosmic expansion than the $\Lambda$CDM model.
In summary, a simple mapping at background level between models in which processes out of equilibrium in the cosmic expansion at late times are allowed and minimally extended versions of the $\Lambda$CDM model can provide a natural and simple way to obtain reinterpretations of the cosmological concordance model in order to provide a more adequate description of the observable universe without invoke the existence of exotic components.
@@ IV Final Remarks
SNaI from Union 2.1 BAO SDSS Final BAO BOSS Fs

### v2

@@ II.3 Modified $\Lambda$CDM model
Notice that the first term in the previous expression resembles the Eckart model, $\Pi=-3H_{0}E_{\Lambda}(z)\xi(z)$ with viscosity coefficient, $\xi(z)=H_{0}E_{\Lambda}(z)$ eckart . However, the above result indicates that we are dealing with a more general description for viscous effects, and our construction does not depend on the usual and intricate constitutive equations which describe such effects; these equations depend on the framework used to introduce the viscosity effects in the cosmic fluid. According to the form of Eq. (24) our construction is more general than the one considered in velten , where the Eckart model was proved. It is worth mentioning that in this case the viscous pressure is reconstructed and characterized by the physical parameters $\alpha$, $\Omega_{i}$ and $H_{0}$ coming from the decaying dark matter model instead of the usual arbitrary parameters $\xi_{0}$ and $s$. In general, the election of the values for the parameters $\xi_{0}$ and $s$ is only justified and motivated by the mathematical simplification that can be obtained on the dynamical equation obeyed by the Hubble parameter in the viscous description.
@@
In Fig. (2) we show the behavior of the viscous pressure (24) as a function of the red-shift for different values of the decaying rate. In any case, we observe that $\Pi$ suffers a transition to an attractor behavior around $z\sim 0.25$. After this point (for $z\leq 0.25$) and throughout the future cosmic history, $\Pi$ evolves exactly the same way, independently of the decay rate value $\alpha$. At the far future, the viscosity pressure in the phantom scenario tends to become more negative as the universe evolves, whilst for the quintessence case it remains negative but close to zero. At times before the attractor transition, the behavior depends on the value of $\alpha$. Namely, for very small $\alpha$ values, $\Pi$ takes large values in the far past and decreases along the cosmic evolution for quintessence and phantom dark energy until today. In contrast, for larger values of the decaying rate ($\alpha>0.1$) in both cases, $\Pi$ increases from negative values in the distant past, towards positive ones. For the phantom case, this transition occurs before than for the quintessence case. For $0.1<\alpha\leq 0.6$ the positive phase of $\Pi$ in the past lasts a short while and finishes around the attractor transition. In the extreme case with $\alpha>0.8$ this phase does not occur. In addition, we computed the deceleration parameter (see figure (3)) and observe that it exhibits a transition from positive to negative before the change of sign in $\Pi$. On the other hand, as the value of the parameter $\alpha$ increases, such a transition for $q$ is lost, and the deceleration parameter is always negative. This is interesting as the viscous effects give rise to an accelerated expansion even for the quintessence case, in which the state parameter of dark energy is not as negative as for a CC. Notice that since in our formulation we have $\Lambda\rightarrow\Lambda(z)$, the behavior of $q$ at late times deviates from the expected value $-1$ obtained for the CC due to the existence of the viscous contribution in the matter sector. However, as can be seen in the plot, such an attractor can be recovered by considering the appropriate value for $\omega_{\mathrm{de}}$.
@@
Finally, using Eq. (8) with $p=0$ we can construct the $\Gamma$ term if we interpret $\Pi$ as the creation pressure, in this case we have
@@
where $\Omega_{\mathrm{vm}}(z)$ is obtained from Eq. (22) and $\Pi$ can be taken from (24). The source term obtained in this description deviates from the usual functional form used in the literature for $\Gamma$ gamma , that is, $\Gamma\propto E^{m}(z)$, being $m$ a constant. In this case we observe, $\Gamma=\Gamma(E,E^{\prime})$, where the prime denotes derivatives w.r.t. red-shift. This behavior is interesting since incorporates matter couplings with curvature, see, for instance, Ref. odintsov . Given the form of $\Pi$, we observe that the particle production behaves in a different way for quintessence and phantom dark energy and will be well defined always that $\Gamma>0$ since the particle production is possible due to the transfer of energy from the gravitational field to matter creation .
@@
III Viability of the model based on observations
III.1 Test of the expansion of the universe with Supernovae IA from Union 2.2
In this section we aim to test the capability of our model to predict the accelerated expansion of the universe. Firstly, in order to make a minimal test, we compare our theoretical predictions of luminous distance modulus as a function of redshift with measurements from the sample of supernovae type IA from the Supernovae Cosmology Project compilation Union 2.2 Suzuki:2012 . It is important to mention that in this work we limit ourselves to considering a small baseline sample of supernovae given that our purpose is to make a minimal test of the viability of our model at the level of the background cosmology and to obtain rough estimates for the parameters rather than determine precise Bayesian inferences of their values. Once this minimal test is made, in further work we aim to extend our analysis and strengthen the results presented here by considering, on one hand larger and more recent datasets such as Pantheon, DES, LIGO etc. PantheonScolnic:2021 ; DES:2024 ; LIGOScientific:2017 . On the other hand, by taking into account the test for the perturbations, it is worth mentioning that the marginalized distributions obtained here would be useful as priors to such a further analysis. In regard of the background cosmology, considering distance measurements of larger samples of supernovae at larger redshifts such as those of DES and Pantheon, constraints on $\alpha$ are expected to be stronger since the sensitivity of $\mu$ increases at high redshift. Including such data sets would be helpful to tighten constraints on the Hubble constant. Furthermore, comparing theoretical predictions for perturbations within this model with observations of LSS from different surveys would greatly enrich the analysis. For instance, effects of the viscosity on expansion of the universe could be distinguished from those on the structure formation process, specially those on small scales hu2 ; Lopez:2022 , and this would be useful to investigate the scale dependence of the $\alpha$ parameter. In order to obtain the theoretical prediction, we firstly computed the luminous distance as function of red-shift given as:
@@ III Viability of the model based on observations
Where $c$ is the speed of light and the denominator in the integrand corresponds to that given by (23). We carried out the corresponding numerical integrations by implementing the trapezoid method in our own Python code. Since the truncation error of this method is of order 2 of the step size, the accuracy of our results is good enough for our purposes. Afterwards we computed the distance modulus simply by using:
@@
Figure 4: Predicted distance modulus as function of red-shift from our model (purple line) overlapped to the observed modulus for Union 2.2 supernovae. The values of the parameters for this model corresponds to $\Omega_{m}h^{2}=0.126$, $H_{0}=73\,\mathrm{km/s/Mpc}$, $\omega_{0}=-0.7$ and $\alpha=0.5$.
As shown in Figure (4), as the decay rate parameter $\alpha$ increases, a larger distance modulus is expected and this modification is more pronounced at large red-shift values, while predictions for small red-shifts remain indistinguishable from those of $\Lambda$CDM. Given the small sensitivity of $\mu$ to variations of $\alpha$, it is expected that these data cannot strongly constrain this parameter.
@@ III.2 Test of the expansion of the universe with BAO
Given that SNIa distance measurements are not sensitive to variations of $\alpha$, constraints for the parameters are expected to be weak and that degeneration between them exists, therefore, in order to overcome such inconveniences, we additionally consider in our analysis that measurements of the BAO feature along the line of sight and transverse directions can separately measure $H(z)$ and the comoving angular diameter distance $D_{M}(z)$ from SDSS-III data Alam:2017 . Variations in the cosmological parameters or the pre-recombination energy density can alter the sound horizon of acoustic oscillations given by:
@@
where $\mathcal{R}=\frac{3\rho_{b}(z)}{4\rho_{\gamma}(z)}=\frac{3\Omega_{b0}}{4\Omega_{\gamma 0}}(z+1)^{-1}$ and $z^{*}$ corresponds to the red-shift at recombination. Therefore, BAO measurements actually constrain the combinations of $D_{M}(z)/r_{s}$ and $H(z)r_{s}$, therefore we compare the predictions of $D_{M}(z)r_{s,fid}/r_{s}$ and $H(z)r_{s}/r_{s,fid}$ with their inferred values from the observations of BAO by BOSS-DR12, where $r_{s,fid}$ corresponds to the fiducial value of $\Lambda$CDM for the sound horizon given as $r_{s,fid}=147.78\,\mathrm{Mpc}$ which is used as normalization constant. In Figure (5) the predictions of $D_{M}(z)r_{s,fid}/r_{s}$ are shown for different models corresponding to a range of values of $\alpha$. Clearly, this quantity barely depends on $\alpha$ at low red-shift, on the contrary, it is sensitive to this parameter at high red-shift. Within this range $\alpha\sim 0.5$ seems to fit properly the observational data. On the other hand, an angle-averaged galaxy BAO measurement constrains the following combination Alam:2017 :
@@ III.3 Bayesian Analysis
In order to obtain baseline estimates for the parameters of the viscous $\Lambda$CDM model, we sample the corresponding parameter space using the Markov Chain Monte Carlo (MCMC) method. Specifically, we implemented the Metropolis-Hasting algorithm in our own Python code to construct a set of Markov chains that serve as samples in the space of parameters of our model used to determine the posterior distribution function corresponding to supernovae data from Union 2.2 and BAO data from BOSS-DR12 Suzuki:2012 ; Alam:2017 . Since we aim to obtain base estimations of the parameters of this model, we use flat priors within ranges that give rise to a physically admissible background evolution. Specifically, flat priors for parameters within the following ranges are considered: $\Omega_{m}h^{2}\,\in,\ (0,1)$, $\omega_{0},\ \in,\ (-2,0)$, $H_{0},\ \in,\ (30,120)$, $\alpha,\ \in,\ (-2,2)$. Notice that we consider negative values for $\alpha$ which correspond to the case in which dark matter is produced by the inverse decay process; although positive values give rise to more realistic model, we take these cases into account for completeness and because the posteriors are independent of prior choice as desired. We marginalized the posterior distribution to estimate the best-fit parameters of our model and their uncertainty at the level of confidence $68\%$ and $95\%$ (see Figure (6) and the Table (1) at the end of this section).
@@
Interestingly, the marginalized distribution for $\alpha$ is pretty similar according to SNIa and BAO, with a best-fit value $\alpha_{bf}\simeq 0.5$. The fact that $\alpha\neq 0$ gives rise to a better fit of both datasets suggests that the occurrence of dissipative effects along the expansion history is not only feasible, but rather the dynamics of expansion might be better described in comparison to the standard $\Lambda$CDM. Another point that is important to stress is that marginalized distributions for $\Omega_{m}h^{2}$ and $H_{0}$ are consistent even at $1-\sigma$ for both datasets, SNIa and BAO. A slight tension can be observed for the dark energy state parameter $\omega_{0}$. Although this discrepancy in the estimate of $\omega_{0}$ for both datasets is not statistically significant, it can be due to the different features of the expansion of the universe at late and early times. Also, it is worth noticing that estimates for these parameters according to BAO are stronger than those for SNIa, this can be explained by the large sensitivity of the theoretical function $D_{M}(z)$ to these parameters at the largest red-shift values in the sample and to the smallness of the error bars of the BAO measurements (see Figure (5)). It is also worth mentioning that the constrained values for the parameter $\omega_{0}$ are in good agreement with the bounds obtained by the DESI Collaboration for a dynamical dark energy model labeled as $\omega_{0}\omega_{a}$ desi .
To end this section we provide some comments on the consideration of the joint data in our analysis. For the flat priors mentioned above, we obtain the following best-fit values for the model parameters at the $68\%$ confidence level, $\Omega_{m}h^{2}$: $0.1758^{+0.0042}_{-0.0012}$, $H_{0}$: $77.83^{+0.34}_{-0.30}$, $\omega_{0}$: $-1.197^{+0.077}_{-0.077}$ and $\alpha$: $0.30^{+0.16}_{-0.21}$. As can be seen, the value of $H_{0}$ tends to high values in order to satisfy the strong constraint imposed by BAO on $\Omega_{m}h^{2}$. However, this tension is alleviated if the Gaussian prior, $\Omega_{m}h^{2}=0.1434^{+0.0020}_{+0.0020}$, obtained from the constraint of various CMB spectra at $68\,\%$ of confidence level is used in our analysis; see references given in planck1518 for the Planck Collaboration results. The best-fit values in the latter case are also shown in Table (1). We highlight the fact that consideration of joint data together with the Gaussian prior for $\Omega_{m}h^{2}$ provides a better estimation of the parameter $\alpha$ w.r.t. SNIa and BAO datasets since the corresponding uncertainties decrease, and we also observe that the Hubble constant value is closer to that reported by the SH0ES Collaboration shoes . The consideration of more datasets could improve the parameter estimations presented here.
@@
SNaI from Union 2.2 BAO SDSS Final BAO BOSS Fs Joint data
@@
68% limits
$0.1240^{+0.0010}_{-0.00049}$
$75.85^{+0.27}_{-0.24}$
$-0.889^{+0.035}_{-0.040}$
$0.150^{+0.058}_{-0.090}$
Table 1: Values for the best-fit parameters of the model at the level of confidence $68\%$ and $95\%$ and $68\%$ for the joint data analysis.
IV Final Remarks
In this work, we study a cosmological model in which a generic candidate of dark matter undergoes an exponential decay into relativistic dark particles at the late stages of evolution of the universe. Since the decay process of dark matter occurs out of chemical equilibrium given that the average kinetic energy is lower than the mass of the candidate, the entropy grows considerably and therefore it is irreversible. As a consequence, the fluid turns out to be not perfect, and a bulk viscosity pressure can be associated to it.
On the other hand, it is well known that different cosmologies describe identically the background evolution of the universe, that is, the solutions for the scale factor are degenerate between models. Therefore, by virtue of this degeneracy at the background level, a mapping between the space of parameters of different models can be established. In this work, we map the parameters of the cosmological model including decaying dark matter described above and a slight extension of the standard cosmological model in which dark matter corresponds to an effective viscous fluid and dark energy corresponding to a perfect fluid with barotropic equation of state. By setting this mapping, we are able to compute an estimate of the bulk viscosity of decaying dark matter in an effective manner and to describe the expansion of the universe according to this framework. As mentioned in the work, we restricted ourselves to background analysis. Given that the matter sector evolves in time in a different way in both models, some differences such as distinct moments of the matter-radiation equality ($z_{\mathrm{eq}}$) and the time of the last scattering ($z_{\mathrm{LSS}}$) are expected between the models. However, due to the degeneracy condition, the matter density could be adjusted in order to recover the same values for $z_{\mathrm{eq}}$ and $z_{\mathrm{LSS}}$ in the degenerate models. Therefore, the computation of spectrum of anisotropies in the cosmic microwave background and the matter power spectrum could exhibit a breaking of the degeneracy explored in this work at the background level. We leave this subject open for future investigation. As discussed previously, the reconstructed viscous pressure evolves in a different way depending on the value of the parameter $\omega_{\mathrm{de}}$. However, for phantom dark energy and quintessence, it has a negative contribution to the pressure of the dark matter sector from the recent past ($z\leq 1$) until the far future ($z=-1$). If we consider the same region for the deceleration parameter, we can observe a transition from decelerated to accelerated cosmic expansion for $\alpha\leq 0.6$. The behavior at the far future for this parameter deviates from the one expected in the CC scenario since in our case due to the mapping between models, we must reinterpret the CC as a variable dark energy. It is worthy to mention that despite the consideration of the phantom dark energy, $\omega_{\mathrm{de}}<-1$; the model does not exhibit a true (big rip) singularity; in this case the singularity is allowed until infinite time has elapsed; this is known as little rip and it has been explored previously within viscous cosmologies.
In addition, by using Bayesian statistics, we test the expansion of the universe predicted by our model against observations of luminous distance and red-shift of type IA supernovae from the Union 2.2 sample and distance measurements from BAO detected in SDSS catalogs. By sampling the space of parameters of our model using the Markov-Chain-Monte-Carlo method implemented by means of the Metropolis-Hastings algorithm in our own Python code, we derive the corresponding 1D and 2D posterior distributions from which we obtain estimates of the model parameters and their uncertainties. As the main result of this analysis we found that dissipative effects along the cosmic history are feasible and could provide a better description of the cosmic expansion than the $\Lambda$CDM model.
In summary, a simple mapping at the background level between models in which processes out of equilibrium in the cosmic expansion at late times are allowed and minimally extended versions of the $\Lambda$CDM model can provide a natural and simple way to obtain reinterpretations of the cosmological concordance model in order to provide a more adequate description of the observable universe without invoke the existence of exotic components.